Hi i want to know how you get 18 from this word problem. The problem is The height of the flagpole is three fourths the height of the school. The difference in their heights is 4.5m. What is the height of the school? Choices for school's height:18,20,24
Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

4.5 = 1/4 the heigth
so 4*4.5=18

Answer by Alan3354(69443)

Hi i want to know how you get 18 from this word problem. The problem is The height of the flagpole is three fourths the height of the school. The difference in their heights is 4.5m. What is the height of the school? Choices for school's height:18,20,24
----------------------
p = height of the pole, s = height of the school
p = (3/4)s
s-p = 4.5 --> s = p+4.5
p = (3/4)*(p+4.5)
4p = 3(p+4.5) = 3p + 13.5
p = 13.5 = 3s/4
----------------
s = 18
--------------
The easy way:
The difference is 1/4 of the height of the school, and that is 4.5 meters
4*4.5 = 18